OpenStudy (anonymous):

Pleasee Helpp? D: Jessica has been tracking the population of a town. The population was 8,230 last year and has been increasing by 1.25 times each year. How can Jessica develop a function to model this? f(x) = (8230 • 1.25)x where 1.25 is the rate of growth f(x) = 8230(1.25)x where 1.25 is the rate of growth f(x) = 1.25(8230)x where 8230 is the rate of growth f(x) = 8230(x)1.25 where 8230 is the rate of growth

OpenStudy (anonymous):

f(x) = 8230(1.25)x where 1.25 is the rate of growth im positive

OpenStudy (anonymous):

can you explain how to get the answer?

OpenStudy (anonymous):

@hartnn can you explain for me?

OpenStudy (mathmale):

Apologies! This is an exponential growth situation, not a linear one. Assume that last year was Year 0. Then the initial value of the population was 8230, period. Now, some questions for you: Please evalute 1.25^0, 1.25^1 and 1.25^2. Then we'll get back to our discussion.

OpenStudy (mathmale):

Please note that 1.25^2 signifies \[1.25^{2}\] and that if you choose not to use Equation Editor, please write that as 1.25^2.

OpenStudy (mathmale):

@Jessicaa : no further interest in solving this problem?

OpenStudy (anonymous):

1.25^0=1 1.25^1=1.25 1.25^2=1.5625 are my answers

OpenStudy (mathmale):

thanks. that's absolutely right. Notice that every value of 1.25^x is obtained by multiplying the previous one by 1.25. Does this make sense? You've typed 1.25^2=1.5625, which is correct; the next power of 1.25 is 1.5625*1.25. Right? This is why I label this problem an "exponential model."

OpenStudy (anonymous):

Yes I understand what you're saying now

OpenStudy (mathmale):

And note that 1.5625*1.25 is the same as 1.25^3. So, your initial value for the population is 8230 and your common multiple is 1.25, and your exponent is simply x, the number of years in question. You set x= {0, 1, 2, 3, 4, ... }. So, which of the answer choices is the correct one to describe what we've discussed here?

OpenStudy (anonymous):

f(x) = 8230(1.25)^x where 1.25 is the rate of growth right?

OpenStudy (mathmale):

Be really careful about typing 1.25x here; that means "1.25 times x," which is inappropriate in this problem. That's why I emphasized writing this expression as 1.25^x. " ^ " signifies exponentiation.

OpenStudy (mathmale):

Yes, that's great. Looks like you've really come to understand this problem well.

OpenStudy (anonymous):

Yes. Thank you so much!

OpenStudy (mathmale):

My great pleasure! See you again.